# Situationally-aware Path Planning Exploiting 3D Scene Graphs

### v1

@@ IV-B Results and Discussions
Sequential execution results Table II summarizes the efficiency gains of the evaluated methods across all queries. Specifically, in the planning scenarios, S-Path(s) demonstrates multifold gains in efficiency, with an average increase of $5.7\times$. This is even without considering its pairing with RRT*, where in many cases S-Path(s) converges while other ablations do not. The efficiency gains from pairing with PRM* are greater ($8.1\times$ on average) than those from pairing with BIT* ($4.3\times$ on average), because the uninformed PRM* benefits more from the C-space restrictions provided by S-Path(s). Moreover, while S-Path(s) on average produces slightly longer paths (difference in $l_{95\%}$ between $5.4\%$ and $15.9\%$, with an average of $1.9\%$), due to its constraint of passing through the center of doorways; this is compensated for by substantial improvements in the time-to-plan (reduction in $ttp_{95\%}$ between $1.1\times$ and $27.2\times$, with an average of $5.7\times$).
In replanning scenarios, S-Path consistently produces shorter paths, averaging $14.9\%$ shorter than those generated by I. Its heuristics prove especially beneficial in environments with blockages, where increased complexity frequently causes traditional heuristic-guided planners to fail. Furthermore, the ability to cache and reuse previously solved subproblems eliminates the need for redundant computations, further reducing ttpttp and improving overall efficiency, by up to $52\times$ for BIT* and $22\times$ for PRM*. A notable example is the SF2-RS2 with BIT*, where the replanning task becomes nearly trivial, as most subproblems are either cached from the initial planning attempt or are reduced to simple point-to-point queries without obstacles, that are trivial for BIT*.
@@
Figure 8: Parallel execution speedups relative to sequential execution as a function of the number of identified subproblems, evaluated across queries in different environments. Each marker represents a query in the respective environment. On average, speedup is proportional to the number of subproblems, with queries in synthetic environments exhibiting higher speedups due to the absence of irregular obstacles.
Parallel execution results S-Path yields parallel execution speedups that correlate with the number of subproblems, as shown in Figure 8, though the actual gains depend heavily on the subproblem setup, since the subproblems vary in complexity. When sufficient cores are available to process all subproblems concurrently, the total parallel execution time approaches the ttpttp of the largest subproblem. This results in greater variation in speedups for queries with more subproblems, as emphasized by the taller boxplots in Figure 8. Conversely, when there are fewer cores than subproblems, the speedup gains diminish accordingly. However, in all queries, including those in large environments such as SF2 (see area in Table I), the number of subproblems was at most 10, so the number of available cores was not a limiting factor. This is largely due to subproblem merging, which avoids the creation of very small, practically inefficient subproblems (as discussed in Section III-D), making S-Path highly scalable. Furthermore, it is important to note that the semantic planning for all queries took less than $1\,\mathrm{ms}$, making its overhead negligible.
@@ V Conclusion
In this work, we proposed S-Path, a novel metric-semantic path planning algorithm that augments sampling-based planners with semantic information derived from indoor 3D scene graphs, which improves planning efficiency while providing human-interpretable path planning. Experimental results demonstrate that S-Path achieves up to $25\times$ efficiency improvement over the underlying sampling-based planner. This originates from the decomposition of the global planning problem into multiple independent subproblems by exploiting room and doorway semantics, effectively reducing the configuration space. This decomposition also enables parallel execution, allowing for further speedup which correlate with the number of subproblems. In scenarios where the original path becomes infeasible due to blockages, S-Path can efficiently replan by reusing previously solved subproblems, achieving up to $52\times$ overall efficiency improvement. However, S-Path enforces traversal through the center of doorways, which may lead to suboptimal paths when considering wide doorways. Additionally, S-Path currently assumes rooms delimited by planar walls, which might arise challenges in contour generation and sample space restriction in environments with complex curved architectural geometries. Future work will focus on refining path optimality through a local-to-global path optimization step and on extending contour generation to handle more complex architectural structures for improved generalizability.

### v2

@@ IV-B Results and Discussions
Sequential execution results Table I summarizes the efficiency gains of the evaluated methods across all queries. Specifically, in the planning scenarios, S-Path(s) demonstrates multifold gains in efficiency, with an average increase of $6.0\times$. This is even without considering its pairing with RRT*, where in many cases S-Path(s) converges while other ablations do not. The efficiency gains from pairing with PRM* are greater ($8.7\times$ on average) than those from pairing with BIT* ($4.5\times$ on average), because the uninformed PRM* benefits more from the C-space restrictions provided by S-Path(s). Moreover, while S-Path(s) on average produces slightly longer paths (difference in $l_{95\%}$ between $5.0\%$ and $15.9\%$, with an average of $2.5\%$), due to its constraint of passing through the center of doorways; this is compensated for by substantial improvements in the time-to-plan (reduction in $ttp_{95\%}$ between $1.1\times$ and $27.2\times$, with an average of $6.0\times$).
In replanning scenarios, S-Path consistently produces shorter paths, averaging $14.9\%$ shorter than those generated by I. Its heuristics prove especially beneficial in environments with blockages, where increased complexity frequently causes traditional heuristic-guided planners to fail. Furthermore, the ability to cache and reuse previously solved subproblems eliminates the need for redundant computations, further reducing ttpttp and improving overall efficiency, by up to $52\times$ for BIT* and $22\times$ for PRM*. A notable example is the SF2-RS2 with BIT*, where the task becomes nearly trivial, as most subproblems are either cached from the initial planning attempt or are reduced to simple point-to-point queries without obstacles.
@@
Figure 8: Parallel execution speedups versus the number of identified subproblems, evaluated across different environments. Each marker corresponds to a single query. Overall, the speedup increases approximately proportionally with the number of subproblems, with synthetic environments showing higher speedups due to the absence of irregular obstacles.
Parallel execution results S-Path yields parallel execution speedups that correlate with the number of subproblems, as shown in Figure 8, though the actual gains depend heavily on the subproblem setup, since the subproblems vary in complexity. When sufficient cores are available to process all subproblems concurrently, the total parallel execution time approaches the ttpttp of the largest subproblem. This results in greater variation in speedups for queries with more subproblems, as emphasized by the taller boxplots in Figure 8. Conversely, when there are fewer cores than subproblems, the speedup gains diminish accordingly. However, in all queries, including those in large environments such as SF2, the number of subproblems was at most 10, so the number of available cores was not a limiting factor. This is largely due to subproblem merging, which avoids the creation of very small, practically inefficient subproblems (as discussed in Section III-D), making S-Path highly scalable. Furthermore, it is important to note that the semantic planning for all queries took less than $1\,\mathrm{ms}$, making its overhead negligible.
@@ V Conclusion
In this work, we proposed S-Path, a novel metric-semantic path planning algorithm that augments sampling-based planners with semantic information derived from indoor 3D scene graphs, which improves planning efficiency while providing human-interpretable path planning. Experimental results demonstrate that S-Path achieves up to $25\times$ efficiency improvement over the underlying sampling-based planner. This originates from the decomposition of the global planning problem into multiple independent subproblems by exploiting room and doorway semantics, effectively reducing the configuration space. This decomposition also enables parallel execution, allowing for further speedup which correlate with the number of subproblems. S-Path can also efficiently replan by reusing previously solved subproblems, achieving up to $52\times$ overall efficiency improvement. However, S-Path enforces traversal through the center of doorways, which may lead to suboptimal paths when considering wide doorways. Additionally, S-Path currently assumes rooms delimited by planar walls, which might arise challenges in contour generation and sample space restriction in environments with complex curved architectural geometries. Moreover, S-Path assumes an accurate map of the environment initially provided and does not account for uncertainties. Future work will focus on refining path optimality through a local-to-global path optimization step, extending contour generation to handle more complex architectural structures, and incorporating robustness to uncertainties in the environment.